A 5 kg of ice cube at -10°C is mixed with 0.1 kg of water at 80 °C. There is no heat loss to the surrounding. The specific heat capacity of ice is 2.22 kJK-'kg-1,The specific heat capacity of water is 4.187 kJK-'kg-! . The specific latent heat of fusion of ice is 333 kJ kg-1. (a) What is the final physical state of the mixture?


1
Expert's answer
2021-11-30T17:45:01-0500

"c_1m_1(t_2-t_1)=\lambda m_2+c_2m_2\Delta t+c_1m_2(0-t_2)"


"2220\cdot5\cdot(t_2-(-10))=333000\cdot 0.1+4187\cdot 0.1\cdot80+2220\cdot 0.1\cdot (0-t_2)\to"


"t_2=-3.9\u00b0C" , the final physical state of the mixture is ice. Answer